OpenStudy (anonymous):

can somebody please help me with 64 = 8 ( r + 2) ?

OpenStudy (anonymous):

okay so distribute 8(r+2)

OpenStudy (anonymous):

you would get 8r + 16

OpenStudy (anonymous):

remember the distributive property: a(b + c) = (a * b) + (a * c)

OpenStudy (anonymous):

so now we have 64 = 8r + 16

OpenStudy (anonymous):

64= 8r + 16 -16 =8r -16

OpenStudy (anonymous):

48= 8r

OpenStudy (anonymous):

now subtract 16 from both sides 64 = 8r + 16 -16 -16 -------------- 48 = 8r now divide 8 from both sides /8 /8 ---------- r = 6

OpenStudy (anonymous):

r=6

OpenStudy (anonymous):

thank you for the answers